OpenStudy (anonymous):

The rectangle shown has a length (l) of 9 and a width (w) of 7.5. A second rectangle has a length of l + 2.5 and a width of w – 2. Find the area of the second rectangle. A. 33 square units B. 34 square units C. 63.25 square units D. 67.5 square units

OpenStudy (anonymous):

11.5*5.5=63.25

OpenStudy (anonymous):

Using the values for L and w you were given: l=9, w=7.5 Just put them into the second rectangle's equations. l+2.5=9+2.5=11.5 w-2=7.5-2=5.5 so the area is 11.5 * 5.5 = 63.25

OpenStudy (anonymous):

6*5*7=210